To solve the radical equation below, which step would result?
Square root of (x-5) - 3 = 4
Thanks~

Answer by dolly(163)

given sqrt(x-5) - 3 = 4
adding 3 to both the sides,
==> sqrt(x-5) = 7
squaring both the sides,
==> x-5 = 49
==> x = 49 + 5
==> x = 54